OpenStudy (anonymous):

Help please! I keep getting 1 as the answer... Prove the given identity. (sinθ/cscθ) + (cosθ/secθ) = sinθcscθ

OpenStudy (campbell_st):

ok.... well sin/csc = sin / (1/sin) = sin^2 and cos/sec = cos/(1/cos) = cos^2 and sincsc = sin * 1/sin = 1 so you get the base identity sin^2 + cos^2 = 1

OpenStudy (anonymous):

That's what I keep getting. However, the book is asking for me to get sinθcscθ. Of course that just equals 1 but I need to show it as sinθcscθ somehow

OpenStudy (campbell_st):

ok.... weird question sorry I'm not much help

OpenStudy (anonymous):

It's alright. I think I'm just going to solve it regularly and put 1 = sinθcscθ. Thank you though!